### v2

# Zeno Effect Suppression of Gauge Drift in Quantum Simulations

Abstract
Quantum simulation of lattice gauge theories is a promising tool for the study of many complicated problems including ones with real-time dynamics. For gauge theories, however, there is a major challenge in maintaining gauge invariance during time evolution. Such theories have a full Hilbert space that is larger than the physical space—the set of states which are gauge invariant or equivalently respect the Gauss law. While an exact implementation of Hamiltonian dynamics starting in the physical Hilbert space will keep the system in the physical space, various types of errors will inevitably produce components outside of it. This work proposes a method of suppressing this gauge drift via the Zeno effect. As in the standard picture of the Zeno effect, our method relies on frequent projection onto the physical subspace. Additionally, a technique is discussed to reduce the speed of the gauge drift, which helps to reduce the required frequency of projections. We demonstrate our method on a $\mathbb{Z}_{2}$ gauge theory toy model.
preprint: APS/123-QED
I Introduction
Quantum computers promise to be vital tools for the simulation of many fundamental yet complicated problems [1, 2, 3, 4, 5] from a range of subfields, including particle physics, nuclear physics, and condensed matter. The simulation and study of complex systems often requires much more computing power than classical computers can provide [6, 7]; thus, there is a present need to map these systems onto quantum devices. Analog quantum devices manufacture a system that can imitate, at least in some regime, the desired theory while digital quantum devices map a theory onto an array of qubits and control it via quantum gates [8]. Much research has gone into studying these quantum devices and how they can be used for quantum simulation [9].
Like many complicated theories, lattice gauge theories can, in principle, benefit enormously from efficient and reliable quantum simulation [10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25]. These theories typically place matter degrees of freedom on the sites of the lattice while gauge field operators are placed on the links; it is typical to work in a Hamiltonian formulation with discretized space but continuous time [26, 27, 28]. This is usually done in the temporal gauge [29] (which sets the time component of gauge fields to zero) and by taking the continuous time limit. Note that the temporal gauge does not fully exhaust the gauge freedom, leaving a residual gauge freedom for the spatial components. Transformations of this residual gauge are generated by Gauss law operators (in analogy to classical electrodynamics). Physical states are then defined as those that are invariant under all residual gauge transformations.
Much work has gone into the study of both Abelian and non-Abelian lattice gauge theories. Notably, maintaining non-Abelian gauge invariance is particularly difficult. When it comes to analog quantum simulation, many methods have been developed to address this gauge invariance problem, including constructing a system that makes gauge invaraince a consequence of internal symmetry such as angular momentum [30, 31]. Another analog simulation method is to impose an energy penalty on the unphysical states [32, 33, 16, 34, 35, 36, 37, 38, 39, 40], often by adding a term to the Hamiltonian proportional to the square of the Gauss’ law operators. Then, the lowest energy levels of the system are the physical ones, so keeping the system in the low energy regime results, to good approximation, in keeping the system within the physical subspace of the much larger Hilbert space. Variations on this method have been developed, including one where the addition to the Hamiltonian is directly proportional to the Gauss’ law operators [41]. Another such variation on the energy penalty method uses local pseudogenerators, which exactly match the Gauss’ law operators in the relevant physical subspace but not in unphysical sectors, allowing these operators to be much less complex and thus cheaper to engineer in practice [42, 43, 44]. Interestingly, for sufficiently large energy penalties, the protection term added to the Hamiltonian, constructed with local pseudogenerators, can be seen as a strong projector which constrain the dynamics of the system to quantum Zeno subspaces within a given timescale [44, 45].
The method presented in this paper concerns digital quantum simulation. Ideally, any method to simulate a non-Abelian gauge theory of interest, such as quantum chromodynamics (QCD), would only simulate the gauge invariant (i.e. physical) states. Methods of this kind have been proposed [46, 47], which encode only physical degrees of freedom. While these methods don’t have to deal with unphysical states, they are fundamentally more challenging. Thus, many methods simulate all states in the Hilbert space of the system, and must contend with keeping the system within the exponentially smaller physical subspace. Crucially, work has been done to develop methods of time evolution that respect gauge invariance [13, 48]. Thus, ideally time evolving an initial state that is constructed to be physical subspace would keep the system within the physical subspace. However, quantum noise, gate errors, and various errors from approximations can lead to the system picking up contributions from unphysical states. Previous work has addressed this problem in a variety of ways. Perhaps the most straightforward is to simply accept the unphysical contributions on the premise that these contributions are reduced as the time step is decreased [15]. Another method is dynamical decoupling [49], a concept from quantum control theory where local symmetries are maintained via periodic driving. Also, oracles have been developed [50] to check for violations of Gauss’ law for Abelian lattice gauge theories, with similar oracles for non-Abelian theories under development.
Furthermore, a method of engineering classical noise to constrain the system to the physical subspace [51] has been proposed. This method serves as a classical analog to the quantum Zeno effect by using classical fluctuations of a perturbing field to dissipate unwanted evolution into unphysical regions of the Hilbert space. The method of this paper, in contrast, rests on the standard quantum Zeno effect [52], which uses frequent projections to keep the system within a desired space; for our purposes the desired space is of course the physical subspace of the system’s Hilbert space.
This paper is organized as follows: section II outlines our method, which takes inspiration from the rodeo algorithm[53]—a method of projection—and a technique of unitary gauge drift suppression. Section III demonstrates our method on a toy model of simulating pure $\mathbb{Z}_{2}$ gauge theory on a 2x2 lattice with periodic boundary conditions. Section IV goes into more detail about the implementation of our method, and Section V provides concluding thoughts and discusses the required future work.
II Method
The method exploits the quantum Zeno effect to maintain a close approximation of gauge invariance. For a non-Abelian theory, the gauge invariant, or physical, states are given by those states |$\ket{\psi}$ that satisfy
$G^{a}_{x}\ket{\psi}=0$ (1)
for all $a$ and $x$, where $G^{a}_{x}$ is the $a$th Gauss’ law operator at site $x$.
To leverage the Zeno effect, we must invoke a method of projection onto the physical subspace. Our method adapts previous work on the rodeo algorithm [53, 54, 55, 56]. We also incorporate a method of suppressing coherent gauge drift, discussed below.
II.1 Projection via the Rodeo Algorithm
The Rodeo algorithm was developed to prepare eigenvectors of a given Hamiltonian. See figure 1 for the circuit diagram of the algorithm [53].
Figure 1: The circuit diagram for the rodeo algorithm [53]
To make clear how the rodeo algorithm works, we will write out how it works for one ancillary qubit. Starting with the system in the state |$\ket{\psi_{I}}$ and the ancillary qubit initialized to |$\ket{1}$, we have
$\begin{bmatrix}\frac{I}{\sqrt{2}}&\frac{I}{\sqrt{2}}\\ \frac{I}{\sqrt{2}}&-\frac{I}{\sqrt{2}}\end{bmatrix}\begin{bmatrix}I&0\\ 0&Ie^{iEt}\end{bmatrix}\begin{bmatrix}I&0\\ 0&e^{-iH_{obj}t}\end{bmatrix}\begin{bmatrix}\frac{I}{\sqrt{2}}&\frac{I}{\sqrt{2}}\\ \frac{I}{\sqrt{2}}&-\frac{I}{\sqrt{2}}\end{bmatrix}\begin{bmatrix}0\\ \ket{\psi_{I}}\end{bmatrix}=$ (2)
$\begin{bmatrix}\frac{1}{2}(I-e^{-i(H_{obj}-E)t})\ket{\psi_{i}}\\ \frac{1}{2}(I+e^{-i(H_{obj}-E)t})\ket{\psi_{i}}\end{bmatrix}$ (3)
A successful implementation of the rodeo algorithm measures every ancillary qubit in the state |$\ket{1}$ at the end. In the single ancillary qubit example above, this results in the transformation |$\ket{\psi_{I}}\rightarrow\frac{1}{2}(I+e^{-i(H_{obj}-E)t})\ket{\psi_{i}}$ where $E$ is the energy of the target eigenstate and $H_{obj}$ is the Hamiltonian of the system. Decomposing our initial state into energy eigenstates |$\ket{\psi_{i}}=\sum\limits_{n}c_{n}\ket{n}$, we can see that the probability for each energy eigenstate $n$ transformations from $|c_{n}|^{2}$ before Rodeo projection to $|c_{n}|^{2}|\frac{1}{2}(1+e^{-i(E_{n}-E)t})|^{2}=|c_{n}|^{2}\cos^{2}((E_{obj}-E)\frac{t}{2})$ afterwards. Thus, the probability for the system to be in a particular energy eigenstate $E_{obj}$ is multiplied by a factor $\cos^{2}((E_{obj}-E)\frac{t}{2})$, which is notably 1 if $E=E_{obj}$ and less than 1 otherwise. This is how the rodeo algorithm works: it constructs an approximate version of the state of energy $E$ by suppressing all other states with a $\cos^{2}(\frac{1}{2}t\Delta E)$ factor. Stronger suppression of contributions from unwanted energy eigenstates can be accomplished with multiple rounds of this algorithm, as discussed in [53, 54, 55]. Note that in order to implement a round of the rodeo algorithm, a time $t$ must be chosen. If the times are chosen strategically for the various rounds, remarkably strong suppression can be achieved with very modest resources [56].
We note here that the Rodeo projection will fail to project onto the state of interest with finite probability (i.e. if any ancillary qubit is not measured in |$\ket{1}$), where failure means that the resulting state is orthogonal to the target state. Crucially, we note that the probability of a successful projection is proportional to the overlap of the initial state of the system with the target state of the projection. As we plan to use this projection to leverage the Zeno effect, which means making sufficiently many measurements so that the overlap with the target subspace is close to unity, the rate of failure can be made negligibly small. The chance of failure serves as the probabilistically defined time span for which our method works; for example, if the probability for a failed projection is one in a million, we would expect one failure within one million successive projections.
The rodeo algorithm is designed to construct eigenstates of a system’s Hamiltonian. For our purposes, we use the ideas of this algorithm for a different hermitian operator, that we will call $G^{2}$:
$G^{2}=\sum\limits_{a,x}(G^{a}_{x})^{2}$ (4)
From the above definition of the Gauss’ law operators, we see that the physical states are eigenstates of $G^{2}$ with eigenvalue 0. Furthermore, all the other eigenstates will be unphysical and with eigenvalues greater than 0, since $G^{2}$ is positive definite by construction. It is this spectrum of $G^{2}$ that we will exploit when using the rodeo algorithm: instead of constructing some eigenstate of a given Hamiltonian, we will construct a physical state by constructing a ground state of $G^{2}$ via the Rodeo algorithm. This means replacing $H_{obj}$ in 1 with $G^{2}$ as well as removing the phase rotation gates (since we would be aiming for an eigenstate with an eigenvalue of 0, so the required phase rotation is 0*t = 0).
There is one minor detail that we must clarify before moving on. The effectiveness of this projection is helped if the spectrum of the operator, in our case $G^{2}$, is discrete. That way there will be less contamination from the eigenstates nearest to the target eigenstate. Luckily, the spectrum of $G^{2}$ for our principal case of interest $SU(3)$ gauge theories, is in fact discrete. This can be seen very straightforwardly using basic quantum mechanics, though We will point out one small detail: the basic argument implicitly relies on the fact that all states have non-negative norms. While this is obviously true for physical states it is not necessarily true for the full Hilbert space of QCD. However, we are working within a Hamiltonian formulation and using the temporal gauge, which is a ghostless gauge [57], meaning that we do in fact have non-negative norms for all states. Thus, all components $G^{a}$ have discrete spectra, so $G^{2}$ has a discrete spectrum as well. This means the ground state of $G^{2}$ is gapped, which is helpful when it comes to projection.
II.2 Suppressing Coherent Gauge Drift
The basic idea of the Zeno effect is that if you project back to the desired space before straying too far out of it (i.e. frequently enough), then the system stays in the desired space with high probability. This is for a simple reason: projections depend on the probability that the system is in the desired state, which corresponds to the square of the magnitude of the amplitude. Thus, if one projects when the amplitude for the system drifting into an unwanted state is some small value $\epsilon$, the probability of failure is $\epsilon^{2}$—a much smaller value.
Consider what this means in practice if one wishes to study a system’s time evolution over a time, $T$, and gauge-invariance-violating errors in the amplitude accrue at a rate $r$. Suppose one makes $N$ evenly spaced projections (which for the sake of this analysis are assumed to be perfect) during the time evolution. The probability of a failure—obtaining a state outside the physical space—at each measurement is $(rT/N)^{2}$, and the probability of a successful run for the full time evolution (i.e. the probability of $N$ successes and 00 failures) is given by
$P_{s}=\left(1-\frac{r^{2}T^{2}}{N^{2}}\right)^{N}\approx 1-\frac{r^{2}T^{2}}{N}$ (5)
where the approximate equality holds in the limit of large $N$ and $1-P_{s}\ll 1$. The key point is that for $N\gtrsim r^{2}T^{2}$ there is a substantial probability that the full time evolution can occur with the system restricted to the physical space. Note that this means that by making sufficient measurements, the evolution can stay in the physical space even if the evolution time is much larger than $1/r$. This a major improvement.
Of course, if one makes a finite number of projections there is a finite probability of failure—obtaining a state outside the physical space. If that happens, the only recourse is to stop the calculation and begin again—an expensive prospect. Thus, there is an optimization problem in choosing $N$: if $N$ is too large, there is the computational cost of making more projections than are necessary, whereas if $N$ is too small than the failure rate is too high requiring the additional cost of starting over. The optimal value depends on the computation costs of the projection as well as the costs of running the simulation.
With this in mind, the method has been enlarged to include an additional way to suppress gauge drift, an approach developed in ref. [58]. we show that a mixed approach using both quantum Zeno and the the approach of ref. [58] is likely to be more efficient than either one separately.
The method of ref.  [58] is essentially an alternative version of the energy penalty method for suppressing coherent gauge drift. Here, we define coherent gauge drift as a unitary process $U_{D}$ that transforms a physical state |$\ket{\psi}$ to
$U_{D}\ket{\psi}=\sqrt{1-\epsilon^{2}}\ket{\psi}+\epsilon\ket{\omega}$ (6)
where |$\ket{\omega}$ is a nonphysical state and $\epsilon$ characterizes the strength of the drift. Recall that an energy penalty method constructs a Hamiltonian $H_{G}$ with the behavior that $H_{G}\ket{\psi}=0$ for physical states |$\ket{\psi}$ and $H_{G}\ket{\omega}>0$ for unphysical states |$\ket{\omega}$. Adding $\lambda_{G}H_{G}$ for $\lambda_{G}\gg 0$ to the Hamiltonian of the system suppresses transitions to unphysical states via an energy penalty. The authors of [58] note that in principle, building $H_{G}$ out of the fundamental fields of the lattice gauge theory would be quite difficult for a non-Abelian theory; however, quantum simulation via Trotterized time steps depends upon $e^{-iH_{G}t}$ not solely $H_{G}$. Furthermore, they point out that this operator basically performs a random gauge transformation on the state of the system, which leaves physical states unchanged while unphysical states take on a phase. Thus, this approach dispenses with the difficult task of building $H_{G}$ and instead simply implements the principal effect of including $H_{G}$ via a random gauge transformation. In summary, the method of [58] is to implement, after every Trotterization time step, a random gauge transformation.
Our method aims to leverage this method wholesale: frequent projections back to the physical subspace is expensive, so we include a random gauge transformation after each Trotterization time step. This reduces the gauge drift per time step, allowing us to do projections to the physical subspace less frequently while still staying close enough to the physical subspace to leverage the quantum Zeno effect, which allows the suppression of whatever gauge drift that survives the random gauge transformations.
II.3 Calibration
Our method, as outlined above, is to conduct a random gauge transformation after every Trotterization time step as well as to conduct projections in the form discussed above to project back to the physical subspace. One important detail that needs to be specified is the frequency of projections. To this end, we include a calibration phase in the method—a set of initial studies—to best estimate the frequency of projections that constitutes a useful compromise between the minimization of gauge drift and the minimization of the number of projections. We note up front that in principle doing a projection after every time step would be most effective in terms of reducing gauge drift as the probability of failure of the projection is very low. The problem with this is simply that it is extremely expensive computationally. In the next section we will outline results from a toy model that show that other, less expensive forms of the method work well. This suggests that realistic calculations can balance the resource costs with the requirements on effective gauge drift suppression.
We propose an approach that should be useful in situations when a simulation needs to be done repeatedly—in that case, a simple calibration phase to optimize the calculation is done at the outset: start with an initial trial frequency of projections based on whatever information and intuition one has about the system. Here, any previous knowledge about the system being simulated should be exploited, especially concerning the speed of gauge drift. With an initial frequency chosen, the time evolution of the system is run a small number of times with the chosen trial frequency of projections to see how long before the projection step fails. If it appears that failure will in general occur before the desired length of time of the simulation, the frequency of projections can be increased. This process can be repeated until an acceptable balance is found.
Behind this simple calibration phase is the fact that the average gauge drift is linearly proportional to the number of time steps between projections. Intuitively, this also means the probability of a failed projection is linearly proportional to the number of time steps between projections. With these relations in mind, it is simple to tune the frequency of projections given a relatively few trial runs.
III Performance
To study this we consider a small “toy” model. It is a pure $\mathbb{Z}_{2}$ gauge theory in 2 spatial dimensions [59]. We simulate this on a 2x2 lattice with periodic boundary conditions. The purpose of using such a small simulation is simply that it allows one to explore the method with very little cost. The Hamiltonian for the toy model is
$H=-k\sum\limits_{j}\square_{j}-g\sum\limits_{i}\sigma^{x}_{i}$ (7)
where $\square_{j}$ is the $j$th plaquette operator, $\sigma^{x}_{i}$ is the Pauli x operator on the $i$th link, and $k$ and $g$ are parameters. A plaquette operator in this context is a product of the $\sigma^{z}$ operators living on the four sides of a plaquette in the lattice.
In this system, the Gauss’ law operators are constructed by taking the product of the $\sigma^{x}$ operators living on the four links emanating from a site on the lattice. We note here that while our method in general works by constructing an operator $G^{2}$ consisting of a sum of the square of all Gauss’ law operators, we have to modify this approach for this toy model as the square of a product of $\sigma^{x}$ operators from different links is manifestly the unity operator. Instead, we use a direct sum of the four Gauss’ law operators of this system, which we will call $G_{tot}$. We also note that for this system, we define a physical state as one satisfying
$G_{a}\ket{\psi}=\ket{\psi}$ (8)
for $a=1,2,3,4$ denoting the associated site. As each $G_{a}$ has a 2-level spectrum (1 and -1), we see that $G_{tot}$ has a 3-level spectrum (-4, 0, and 4), with the physical states satisfying $G_{tot}\ket{\psi}=4\ket{\psi}$.
Our aim is to study gauge drift, so we include in our toy model an artificial gauge drift operator $D$ with each time step. We base our constructed gauge drift operators off of eq. (6); specifically, we connect each of the, in our case 32, physical states of the system to a randomly selected unphysical state in exactly the way laid out in eq. (6). All other states are unchanged. This provides a way for the system to drift out of the physical subspace of its Hilbert space in a unitary way, and a parameter, $\epsilon$, to control the strength of this drift. For our simulations, we set this parameter to $\epsilon=0.01$. Note that each time our method calls for time evolution, we select a different set of 32 unphysical states to appear in the gauge drift operator so as to not bias any unphysical states.
With this defined, the algorithm can be used for the toy model. Each time step takes a state |$\ket{\psi}$ to $G_{i}DU_{t}\ket{\psi}$ where $G_{i}$ is a randomly chosen gauge transformation, $D$ is a gauge drift operator, and $U_{t}=\exp{-iHt}$ ) is the standard time evolution operator. The remaining part of the algorithm is the rodeo projection, which occurs after a set number of time steps (determined in the calibration phase). As discussed above, a successful projection takes a state |$\ket{\psi}$ to $\frac{1}{2}(1+\exp(-i(G_{tot}-4)t_{r}))\ket{\psi}$ where here we note that the time used within the projection $t_{r}$ is in general different from the time step size $t$. Recall that all physical states are eigenstates of $G_{tot}$ with an eigenvalue of 4, so they are unaffected while unphysical states will suffer destructive interference.
Figure 2: Comparing gauge drift suppression methods, including no suppression methods (green), a random gauge transformation after each time step (blue), and a random gauge transformation followed by a rodeo projection after each time step (orange). The right plot is the same as the left plot without the green plot so as to zoom in on the behavior of the other two.
Figure 2 demonstrates the basic performance of the method. It shows no suppression, a method of using a random gauge transformation after each time step, and our method of using a random gauge transformation after each time step plus projections, in this case after every time step. It can be seen both that the use of random gauge transformations after every time step greatly suppresses the gauge drift as well as that frequent rodeo projections does in fact keep the system very close to the physical subspace.
Figure 3: Comparing performance of regimes with no gauge transformations and various frequencies of rodeo projections with regimes with gauge transformations and various frequencies of rodeo projections
In the rest of this section we will consider the frequency of projections. To study this, we have run simulations to test many different regimes to compare their effectiveness. In particular, systems without any gauge transformations after time steps with a frequency of rodeo projections ranging from after one time step to after 10 time steps. We also considered systems with a random gauge transformation after each time step with a frequency of gauge projections ranging from after one time step to after 30 time steps. For each regime, we ran our algorithm on 200 systems and calculated the average of the farthest from physical each run reached (which is to say the lowest value of $\expectationvalue{G_{tot}}$).
Figure 3 shows these results. It is clear that including the random gauge transformations into the method significantly helps to reduce the required frequency of gauge projections. In particular, it shows the following approximate equivalences: a projection after every time step performs similarly to a projection every two time steps when gauge transformations are included; a rodeo projection after every two time steps performs similarly to a gauge projection every five time steps when gauge transformations are included; and a gauge projection after every three time steps performs similarly to a gauge projection every ten time steps when gauge transformations are included. This shows that including the gauge transformations reduces the required frequency of rodeo projections by more than half. In particular, we see that for a system that drifts out of the physical subspace slowly, one could perform projections infrequently and still be able to utilize the Zeno effect to great advantage.
IV Implementation
It is important to consider the implementation of this method, as well as their costs. We will start with a conisderation of the implementation of the gauge transformations, which is discussed in [58]. First of all, there is the cost of sampling the gauge transformations, which includes a small number of ancillary qubits; notably, this cost does not scale with the size of the system [48, 58]. When it comes to implementing the sampling, we make use of a common method of sampling uniformly from the Haar measure of a group $G$ wherein a small number (e.g. 2) elements of $G$, call them $g_{1}$ and $g_{2}$, are chosen and then strings (of a chosen length) of these elements are randomly sampled, e.g. $g_{1}g_{2}g_{2}g_{1}g_{2}$. Importantly, this method of sampling elements of $G$ converges to the Haar measure in the limit of long strings. For the present purpose, however, we note that we do not rely on a truly uniform sampling of gauge transformations; thus, relatively short strings can be used, in a similar way to the circuit described in [60].
Next, we consider the implementation of the gauge transformations, which, as the original authors of this method noted in ref.  [58], ”requires about as many G-multiplication gates as the original quantum simulation did”. Thus, the total cost of using gauge transformations in this method is about double what it would be for a straightforward simulation.
Additionally, the implementation and cost of the gauge projections needs to be considered. We have suggested that the rodeo projection method appears to be useful. Note that previous work has shown that the rodeo algorithm can be optimized and made far more efficient [56] via considered choices for the parameter $t_{r}$. We suggest using this optimized approach makes sense for our purposes as well. This work notes that the dominant cost of the rodeo algorithm is the controlled time evolution gate (in our case it is then the controlled U gate with $U=\exp(-iG^{2}t_{r})$); thus, they argue that the total time to conduct a rodeo projection acts as a proxy for computational costs and then proceeds to lay out a method for selecting the $t_{r}$s that maximize efficiency of the algorithm given fixed computational resources [56]. We will also note that constructing $G^{2}$ will be much more difficult than $H$ so our version of a rodeo projection suffers a much heftier cost in construction of the required gates, but so will alternative projection schemes.
Furthermore, the projections themselves require a number of ancillary qubits. Notably, the projections only use one ancillary qubit at a time, so if mid-circuit measurements are allowed then only one ancillary qubit is required, as it can be recycled again and again. If this is not the case, then a significant number of ancillary qubits must be used. Specifically, we estimate one projection would need as many ancillary qubits as sites in the lattice.
V Conclusion
We have outlined an algorithm developed to maintain gauge invariance (or more precisely suppress gauge drift) during the simulation of lattice gauge theories. Our algorithm was based off of the Zeno effect, wherein we use frequent projections (i.e. altered rodeo projections) to keep the system within the physical subspace of its exponentially larger Hilbert space. We also include a technique of performing gauge transformations after every time step to help reduce the system’s gauge drift into the unphysical sector which in turn allows for less frequent projections while still taking full advantage of the Zeno effect. We have shown that our method is effective for a simple small toy model of pure $\mathbb{Z}_{2}$ gauge theory. Future work aims to test our algorithm out on larger systems, and more complicated gauge theories such as $SU(N)$ or gauge theories that include matter.
Acknowledgements.
This work was supported in part by the U.S. Department of Energy, Office of Nuclear Physics under Award Number(s) DE-SC0021143 and DE-FG02- 93ER40762.